UT 20120209 - AO2 commissioning

Observer: SEsposito, EPinna, APuglisi, MXompero, KBoutsia (LBTO)
Telescope Operator: Steve Allanson (LBTO)
Additional AO Support: SEsposito, EPinna, APuglisi, MXompero
Telescope Support: John L.(LBTO)
Instrument Support: Tim S. (LBTO)

Summary

The retroreflector is mounted on the left side during the afternoon, since the sky is completely overcast during the whole day.
Thick clouds are clearing out just before sunset. Try to run some tests on the right side.
After sunset since the clouds seemed to dissipate we decided to take the retroreflector off and go on with AO testing on the left side.
It has been a productive night despite the forecast and for bright stars the AO group obtained >95% strehl ration in H and ~70% in J. For a star of mag 10.5 the strehl obtained in H was ~60%.
Poor seeing for the second half of the night (variable between 1.5'' -2.5'' on the dimm) so some back up tests were executed.

Details

01:50 The telescope is collimated
  • the tests on the right side give error because Pisces is offline

02:01 Sky still looks fairly clear and the AO group decides to remove the retroreflector

02:30 Retroreflector off - steve is opening the dome
  • switch off tree-house ventilation

02:59 Problem with turning on the AGW unit

03:22 iif_preset,'AO1071',GS=3,RS=0,MODE=ACE,PA=180
  • failed because the preset was send only to the left side but on the telescope also the right side is authorized
  • we ask to unauthorize the right side

03:24 iif_preset,'AO1071',GS=3,RS=0,MODE=ACE,PA=180
  • problems with the GCS - it is not resuming guiding - guiding is hung
  • restarting GCS
  • the script exited without error ... (??)

03:29 iif_preset,'AO1071',GS=3,RS=0,MODE=ACE,PA=180
  • problems with the ACE presets ... the argument wfs_name is "none" while when we send the same command from the AOSgui it is set to "flao_wfs"

03:30 iif_preset,'AO1071',GS=3,RS=0,MODE=ACTIVE,PA=180
  • the star was mag=7.2 and they obtained >95% strehl ration in H and ~70% in J.

05:58 trying to take images in open loop we lost the star because the guiding is not working
  • Steve needs to reboot the AZcam and GCS

06:04 iif_preset,'AO1071',GS=3,RS=0,MODE=ACTIVE,PA=180
  • failed - GCS still giving problems
  • repeat

06:06 iif_preset,'AO1071',GS=3,RS=0,MODE=ACTIVE,PA=180
  • preset cancelled by Steve since the GCS is not working ...

06:14 iif_preset,'AO1071',GS=3,RS=0,MODE=ACTIVE,PA=180
  • restart a different server in order to make it work

06:17 iif_preset,'AO305',GS=2,RS=0,MODE=ACTIVE,PA=180
  • changed star because the previous was going low in elevation
  • mag=7.2 also for this star

07:14 iif_preset,'AO327',GS=7,RS=0,MODE=ACE,PA=0
  • this time ace worked
  • iif_runao, retval
  • try some offsets to see if they work
  • iif_offset, 0, 0.5 - test successful
  • for this star with mag 10.5 the measured strehl in H band is ~60%

08:04 iif_preset,'AO860',GS=1,RS=0,MODE=ACE,PA=180
  • mode ace went through again.
  • wfs keyword had been previously set to flao_wfs manually. Probably the script does not set this keyword... and if it finds it ready it goes through, if not it hungs.
  • iif_runao, retval - runao concluded without giving errors but the AO system did not manage to keep the loop closed
  • the loop is open, we send the preset again

08:17 iif_preset,'AO860',GS=1,RS=0,MODE=ACE,PA=180
  • iif_runao, retval

09:12 iif_preset,'AO472',GS=4,RS=0,MODE=ACE,PA=-90
  • iif_runao, retval
  • AO star mag =12.5
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = AcquireRef[260] complete -  dx:-0.388 dy:1.687 slNull: f1:Filter #0 f2:Filter #0 freq:625 mag:12.404 nBins:2 nModes:120 r0:0 snMode:0 strehl:0 ttMod:3
        err = Resmsg[2]  = Safe skip active - cannot proceed (-10045) IDL_SAFE_SKIP_ACTIVE
        err = Resmsg[3]  = StartAO[261] failed - Safe skip active - cannot proceed (-10045) IDL_SAFE_SKIP_ACTIVE
  • Alfio closes the loop manually from the AO side.

09:43 the seeing is >1.5'' on the GCS and ~2.0'' on the dimm which is out of AO specs
  • seeing is getting worse with peaks of 2.7"
  • temperature has dropped 2deg so we probably also have some dome seeing

09:53 iif_preset,'AO420',GS=1,RS=0,MODE=ACE,PA=180
  • we are going to a brighter star to run other tests
  • this one is mag=7.2
  • iif_runao, retval
  • while running some tests, Alfio manually offset the star but the offloads were too high and they lost the star
  • repeat preset

10:09 iif_preset,'AO420',GS=1,RS=0,MODE=ACE,PA=180
  • iif_runao, retval

10:26 going back to the fainter star (12.5mag)
  • iif_preset,'AO472',GS=4,RS=0,MODE=ACE,PA=-90
  • failed because of error with the ccd47
  • repeat
  • iif_runao

10:39 iif_preset,'AO472',GS=4,RS=0,MODE=ACE,PA=-90
  • there is a problem on the AO side ... failed
  • repeat
  • iif_runao, retval
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = Timeout error on baysidey (-5001) TIMEOUT_ERROR
        err = Resmsg[2]  = AcquireRef[334] failed - Timeout error on baysidey (-5001) TIMEOUT_ERROR

10:44 iif_preset,'AO472',GS=4,RS=0,MODE=ACE,PA=-90
  • iif_runao, retval
  • runao successful but right after the loop opened
  • alfio repeats preset from AO giving a more accurate value for the flux

10:53 iif_preset,'AO860',GS=1,RS=0,MODE=ACE,PA=180
  • mag 10.5
  • preset will fail because there are clouds, the background is high and the AO system did not manage to take a dark image...

10:56 iif_preset,'AO860',GS=1,RS=0,MODE=ACE,PA=180
  • iif_runao, retval
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = IDL File error
        err = Resmsg[2]  = AcquireRef[345] failed - IDL File error
* this happened because the estimated star mag is 2 mag fainter because of the clouds and it asked for bin3 but the bin3 table does not exist yet.

11:00 clouds are closing up - we are loosing the star on the guider

11:09 sky almost completetly overcast
  • AO group needs the telescope at the zenith to check something while we are waiting for the clouds to pass

11:44 iif_preset,'AO472',GS=4,RS=0,MODE=ACE,PA=-90
  • iif_runao
  • the same error with the bin tables ... seeing on the dimm 2.2"

12:00 iif_preset,'AO860',GS=1,RS=0,MODE=ACE,PA=180
  • iif_runao
  • the actual mag is 10.4 but the AO system is estimatind ~13
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = Timeout error on baysidey (-5001) TIMEOUT_ERROR
        err = Resmsg[2]  = AcquireRef[352] failed - Timeout error on baysidey (-5001) TIMEOUT_ERROR
* repeat preset and runao
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = AcquireRef[356] complete -  dx:0.468 dy:1.647 slNull: f1:Filter #0 f2:Filter #0 freq:990 mag:11.188 nBins:2 nModes:153 r0:0 snMode:0 strehl:0 ttMod:3
        err = Resmsg[2]  = Safe skip active - cannot proceed (-10045) IDL_SAFE_SKIP_ACTIVE
        err = Resmsg[3]  = StartAO[357] failed - Safe skip active - cannot proceed (-10045) IDL_SAFE_SKIP_ACTIVE

12:32 iif_preset,'HIP58369',GS=4,RS=0,MODE=ACE,PA=180
  • iif_runao
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = AcquireRef[365] complete -  dx:-0.765 dy:0.044 slNull: f1:Filter #0 f2:Filter #0 freq:990 mag:7.696 nBins:1 nModes:400 r0:0 snMode:0 strehl:0 ttMod:3
        err = Resmsg[2]  = Safe skip active - cannot proceed (-10045) IDL_SAFE_SKIP_ACTIVE
        err = Resmsg[3]  = StartAO[366] failed - Safe skip active - cannot proceed (-10045) IDL_SAFE_SKIP_ACTIVE

13:14 iif_preset,'HJ510',GS=4,RS=0,MODE=ACE,PA=0
  • iif_runao
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = Timeout error on baysidey (-5001) TIMEOUT_ERROR
        err = Resmsg[2]  = AcquireRef[423] failed - Timeout error on baysidey (-5001) TIMEOUT_ERROR

13:27 iif_preset,'HJ510',GS=4,RS=0,MODE=ACE,PA=0
  • failed - y slit ...
  • repeat preset - still problems with the slits
  • repeat - this time it was successful
  • iif_runao

13:37 background too high because of twilight - End of Night

-- KonstantinaBoutsia - 09 Feb 2012
Topic revision: r7 - 10 Feb 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback